OpenStudy (anonymous):

A group of students agreed that each would contribute the same amount to buy their favorite teacher a 500.00 php birthday gift. At the last minute, 2 of the students decided not to chip in. This increased the that the remaining students had to pay by 12.50 php per student. How many students actually contributed to the gift?

OpenStudy (ahsome):

Let the number of people be \(x\) \[\text{Number of People}*Pay=Total\]\[x*12.50=500\]\[12.5x=500\]Divide both sides by \(12.5\) \[\frac{12.5x}{12.5}=\frac{500}{12.5}\]Cross the \(12.5\) on the LHS \[\frac{\cancel{12.5}x}{\cancel{12.5}}=\frac{500}{12.5}\] \[x=\frac{500}{12.5}\] Solve for that to find the number of children who payed

OpenStudy (ahsome):

@dabeezy, if you want to clarify ;)

OpenStudy (anonymous):

huh?